# Ramification filtration via deformations, II

### v3

@@ 4.2. Choice of $e\in\mathcal{L}_{O_{M}(\mathcal{K})}$
The structural constants $\eta(\bar{a})\in W_{M}(k)$ reflect the fact that $E$ is a “diagonal element” in ${\mathcal{A}}/J^{p}$ or, equivalently, $E\in\widetilde{\exp}(\mathcal{L}_{O_{M}(\mathcal{K})})$ or, equivalently, $\Delta E\equiv E\otimes E$ modulo $p$-th power of augmentation ideal in $\mathcal{A}^{\otimes 2}_{O_{M}(\mathcal{K})}$. Explicitly, this can be stated via the following condition:
@@
Note that in papers [1, 2, 3] we treated on the first place the simplest choice of such constants $\eta(\bar{a})$ related to the case $e=\sum_{a\in\mathbb{Z}^{0}(p)}t^{-a}D_{a0}$. In this case the constants are quite simple $\eta(\bar{a})=1/r!$, where $\bar{a}=(a_{1},\dots,a_{r})$. However, our approach works equally well also in the above general situation.
@@
There is very interesting possibility to take
@@ 4.3. Statement of the main result
By definition everywhere $D_{a_{i}n_{i}}:=D_{a_{i},n_{i}\mathrm{mod}N_{0}}$, but when computing $\eta(\bar{a},\bar{n})$ or $\gamma(\bar{a},\bar{n})$ it is essential that $n_{i}\in\mathbb{Z}$.
@@
The result about explicit generators of ¯$\bar{\mathcal{L}}^{[v_{0}]}$ can be stated in the following form.
Let ¯$\bar{\mathcal{F}}^{0}_{\gamma,-N}$ be the image of $\mathcal{F}^{0}_{\gamma,-N}$ in ¯$\bar{\mathcal{L}}_{W_{M}(k)}$.
Let $\iota=p^{m}(qp^{M-1}\alpha-(q-1)\beta)\in\mathfrak{A}_{1}^{+}(p)$ be the standard presentation from Section 2.2. Note that by definition $\iota\leqslant p^{M-1}(p-1)b^{*}$, $p^{m}\alpha$ and $p^{m}(q-1)\beta/r^{*}$ are integers not both divisible by $p$. We have also that $p^{m}(q-1)\beta_{\iota}/b^{*}=p^{m+M-1-A}$ where $10\leqslant A\leqslant M-$ 1, compare to Prop. 2.4b). So, by setting $\gamma(\iota):=p^{A}\alpha$ and $M(\iota):=m-A$ we obtain the following presentation of $\iota$:
@@
The condition $\iota\leqslant p^{M-1}(p-1)b^{*}$ holds automatically when we work with elements of $\bar{\mathcal{N}}^{{\dagger}}[0]$. Indeed, if $t^{-\iota}D_{\bar{a}\bar{n}}\in\bar{\mathcal{N}}^{{\dagger}}[0]$ then $-\iota\geqslant-qp^{M-1}|\bar{a}|$. Therefore, if $\iota>p^{M-1}(p-1)b^{*}$ then $|\bar{a}|>(p-1)b^{*}/q$. By condition C3 from Section 2.1 it is equivalent to $|\bar{a}|\geqslant(p-1)v_{0}$ and then by Corollary 1.8, $D_{\bar{a}\bar{n}}\in\bar{\mathcal{L}}(p)_{W_{M}(k)}=0$.
Let $m(\iota)$ be the maximal integer such that $\iota p^{m(\iota)}\leqslant p^{M-1}(p-1)b^{*}$. For any $\iota\in\mathfrak{A}^{+}_{1}(p)$, fix a choice of $m_{\iota}\geqslant m(\iota)$.
@@
¯$\bar{\mathcal{L}}^{[v_{0}]}$ is the minimal ideal in $\bar{\mathcal{L}}$ such that for all $\iota\in\mathfrak{A}^{+}_{1}(p)$ given in notation (4.1), ¯$\bar{\mathcal{F}}^{0}_{\gamma(\iota),-(M(\iota)+m_{\iota})}\in\bar{\mathcal{L}}^{[v_{0}]}_{W_{M}(k)}$.
@@
The elements ¯$\bar{\mathcal{F}}_{\gamma,-n}$ appear as linear combinations of Lie monomials $[D_{\bar{a},\bar{n}}]$ which satisfy the condition $\gamma=\gamma(\bar{a},\bar{n})=a_{1}p^{n_{1}}+\ldots+a_{s}p^{n_{s}}$ with $a_{i}\in\mathbb{Z}^{0}(p)$. There is a lot of monomials containing the generators $D_{0n}$ which do not contribute essentially to the condition $\gamma=\gamma(\bar{a},\bar{n})$. It would be natural to switch from the Lie algebra $\mathcal{L}$ to its quotient $\mathcal{L}^{o}=\mathcal{L}/(D_{0})$ which is related just to the generators $\{D_{an}\ |\ a\in\mathbb{Z}^{+}(p)\}$. Then we obtain the description of the corresponding ramification ideal $\mathcal{L}^{o(v_{0})}$ in terms of generators which all contribute (essentially) to the condition $\gamma=\gamma(\bar{a},\bar{n})$. As a matter of fact, this means that we fixed a lift $\hat{\sigma}$ of the Frobenius automorphism to $\mathcal{G}_{<p,M}$ and work with the Galois group of the totally ramified field extension $\mathcal{K}_{<p,M}^{\langle\hat{\sigma}\rangle}$. This field appears very naturally in the traditional local class field theory as well as in its nilpotent version [23].
@@ 4.4. A recurrent relation
where $\bar{x}_{1}\in\bar{\mathcal{N}}^{{\dagger}}[1,0]$ and $V_{\iota}=\bar{\mathcal{V}}^{0}B^{{\dagger}}(D_{\iota 0}^{{\dagger}})$.
Let $E_{0}=\exp(\sigma^{M-1}\bar{e}_{1}^{(q)})$ and $\exp(\Omega(\sigma^{M-1}\bar{e}_{1}^{(q)})=E_{0}+\mathcal{E}_{0}$. In particular, $E_{0}=\sum_{a\in\mathbb{Z}^{0}(p)}\sigma^{M-1}\eta(\bar{a})t_{1}^{-qp^{M-1}|\bar{a}|}D_{\bar{a},M-1}$ and
@@
Then (4.2) can be rewritten as follows
Note that this relation is still in ¯$\bar{\mathcal{L}}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ due to the following easy lemma:
@@
$\exp(l+l^{\prime})\cdot\exp(-l)-1\in\bar{\mathcal{N}}^{{\dagger}}[1,0]\,.$
As a result, the ideal ¯$\bar{\mathcal{L}}^{[v_{0}]}$ appears now as the minimal ideal in $\bar{\mathcal{L}}$ such that ¯$\bar{\mathcal{L}}^{[v_{0}]}_{W_{M}(k)}$ contains the elements $V_{\iota}\in\bar{\mathcal{L}}_{W_{M}(k)}$ from (4.3).
@@
Let $\iota\in\mathfrak{A}^{0}_{1}$. Set
—  $r(\iota):=\max\{r\ |\ \iota p^{r}\in\mathfrak{A}_{1}^{0}\}(=\max\{r\ |\ |\iota p^{r}|\leqslant p^{M-1}(p-1)b^{*}\})$.
@@
where the last sum is taken over all $\iota\in\mathfrak{A}^{\mathrm{prim}}_{1}$ and $r\geqslant 0$.
Proposition 4.3.
For all $\iota\in\mathfrak{A}^{\mathrm{prim}}_{1}$ and $r\geqslant 0$, ¯$\bar{\mathcal{F}}[ip^{r}]\in\bar{\mathcal{L}}_{W_{M}(k)}$.
The proof will be given in Sections 4.6-4.7 below.
For $r\geqslant 0$, let ¯$\bar{\mathcal{F}}[\iota p^{r}]_{0}:=\bar{\mathcal{F}}[\iota p^{r}]-\sigma\bar{\mathcal{F}}[\iota p^{r-1}]\in\bar{\mathcal{L}}_{W_{M}(k)}$. Then
@@
a) $\bar{x}_{1}\equiv-\sum_{\iota,r}\bar{\mathcal{F}}[\iota p^{r}]t_{1}^{-\iota p^{r}}\,\mathrm{mod}\,\bar{\mathcal{L}}^{[v_{0}]}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}+\bar{\mathcal{N}}^{{\dagger}}(p)$, where the sum is taken over all $\iota\in\mathfrak{A}^{\mathrm{prim}}_{1}$ and $0\leqslant r\leqslant r(\iota)$;
b) if $\iota\in\mathfrak{A}_{1}^{+}(p)$, then
$V_{\iota}\equiv\sigma^{-r(\iota)}\bar{\mathcal{F}}[\iota p^{r(\iota)}]\,\mathrm{mod}\,\left[\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}\right]_{W_{M}(k)}$
@@
Apply induction on $1\leqslant w<p$ to prove the proposition modulo $C_{w}(\bar{\mathcal{L}})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ for all $1\leqslant w\leqslant p$. Clearly, a) holds with $w=1$.
@@
(4.4) $\sum_{\mathrm{ch}\iota=1}t_{1}^{-\iota}V_{\iota}+\sigma{\bar{x}}_{1}^{1}-\bar{x}_{1}^{1}\equiv$
$E_{0}^{-1}\mathcal{E}_{0}-\left(E_{0}^{-1}\sigma\bar{x}_{1}^{1}E_{0}-\sigma\bar{x}^{1}_{1}\right)\,\mathrm{mod}\,[\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}]_{O^{\prime}_{M}(\mathcal{K}^{\prime})}\,.$
Now the RHS belongs to $\bar{\mathcal{N}}^{{\dagger}}[1,0]$ modulo extension of scalars of [¯$[\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}]+C_{w+1}(\bar{\mathcal{L}})$. It is easy to see that the RHS appears in the form
$E_{0}^{-1}\mathcal{E}_{0}+\sum_{\iota,r,a}\bar{\mathcal{F}}[qp^{M-1}a+\iota p^{r+1}]_{0}t_{1}^{-(qp^{M-1}a+\iota p^{r+1})}=\sum_{r\leqslant r(\iota)+1,\iota}\bar{\mathcal{F}}[\iota p^{r}]_{0}t_{1}^{-\iota p^{r}}\,.$
We can assume that in the last sum $r\leqslant r(\iota)$ due to the following Lemma.
Lemma 4.5.
Let $r\geqslant r(\iota)+1$.
a) If $\iota<0$ then ¯$\bar{\mathcal{F}}[\iota p^{r}]t_{1}^{-\iota p^{r}}\in\bar{\mathcal{N}}^{{\dagger}}[0](p)$.
b) If $\iota>0$ then ¯$\bar{\mathcal{F}}[\iota p^{r}]\in\left([\bar{\mathcal{L}}^{{\dagger}[v_{0}]},\bar{\mathcal{L}}]+C_{w+1}(\bar{\mathcal{L}})\right)_{W_{M}(k)}$.
Proof of Lemma.
Part a) follows just from definition of $\bar{\mathcal{N}}^{{\dagger}}(p)$. Part b) provided by the inductive assumption about the RHS. ∎
@@
If $\iota<0$ then the expression for $\bar{x}_{1}$ will be provided with
$\mathcal{S}\left(\sum_{r}\bar{\mathcal{F}}[\iota p^{r}]_{0}t_{1}^{-\iota p^{r}}\right)=-\sum_{j\geqslant 0}\sigma^{j}\left(\sum_{\iota,r}\bar{\mathcal{F}}[\iota p^{r}]_{0}t_{1}^{-\iota p^{r}}\right)=-\sum_{\iota,r}\bar{\mathcal{F}}[\iota p^{r}]t^{-\iota p^{r}}\,.$
Obviously, the elements from $\bar{\mathcal{N}}^{{\dagger}}[0](p)\subset\bar{\mathcal{L}}_{\mathrm{m}^{\prime}}$ will not contribute to the part $\sum_{\iota\in\mathfrak{A}^{+}(p)}t^{-\iota}V_{\iota}$.
@@
$\mathcal{R}\left(\sum_{r\leqslant r(\iota)}\bar{\mathcal{F}}[\iota p^{r}]_{0}t_{1}^{-\iota p^{r}}\right)\equiv\sum_{r\leqslant r(\iota)}\sigma^{-r}\bar{\mathcal{F}}[\iota p^{r}]_{0}t_{1}^{-\iota}\equiv-\sigma^{-r(\iota)}\bar{\mathcal{F}}[\iota p^{r(\iota)}]\,.$
This proves b).
@@
$\mathcal{S}\left(\sum_{r}\bar{\mathcal{F}}[\iota p^{r}]_{0}t_{1}^{-\iota p^{r}}\right)=\sum_{r}\sum_{0\leqslant r_{1}<r}t_{1}^{-\iota p^{r_{1}}}\sigma^{-r+r_{1}}\bar{\mathcal{F}}[\iota p^{r}]_{0}$
$=\sum_{r}t_{1}^{-\iota p^{r}}\sum_{r<r_{1}\leqslant r(\iota)}\sigma^{-r_{1}+r}\bar{\mathcal{F}}[\iota p^{r_{1}}]_{0}=\sum_{r}t_{1}^{-\iota p^{r}}\sum_{\begin{subarray}{c}r_{1}+r_{2}=r\\ r_{2}<0\end{subarray}}\sigma^{r_{2}}\bar{\mathcal{F}}[\iota p^{r_{1}}]_{0}$
$\equiv-\sum_{r}t_{1}^{-\iota p^{r}}\sum_{\begin{subarray}{c}r_{1}+r_{2}=r\\ r_{1},r_{2}\geqslant 0\end{subarray}}\sigma^{r_{2}}\bar{\mathcal{F}}[\iota p^{r_{1}}]_{0}\equiv-\sum_{r}\bar{\mathcal{F}}[\iota p^{r}]t_{1}^{-\iota p^{r}}$
@@
$\sum_{r_{2}+r_{1}=r}\sigma^{r_{2}}\bar{\mathcal{F}}[\iota p^{r_{1}}]_{0}\equiv\sigma^{r}\left(\sigma^{-r(\iota)}V_{\iota}\right)\,\mathrm{mod}\,\left([\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}]+C_{w+1}(\bar{\mathcal{L}}))\right)_{W_{M}(k)}\,.$

### v5

@@ 4.2. Choice of $e\in\mathcal{L}_{O_{M}(\mathcal{K})}$
The structural constants $\eta(\bar{a})\in W_{M}(k)$ reflect the fact that $E$ is a “diagonal element” in ${\mathcal{A}}/J^{p}$ (here $J$ is the augmentation ideal of $\mathcal{A}$). Explicitly, this can be stated via the following condition:
@@
Note that in papers [1, 2, 3] we treated on the first place the simplest choice of such constants $\eta(\bar{a})$ related to the case $e=\sum_{a\in\mathbb{Z}^{0}(p)}t^{-a}D_{a0}$. In this case the constants are quite simple $\eta(\bar{a})=1/r!$, where $\bar{a}=(a_{1},\dots,a_{r})$. However, our approach works equally well in the above general situation.
@@
There is a very interesting possibility to take
@@ 4.3. Statement of the main result
By definition, everywhere $D_{a_{i}n_{i}}:=D_{a_{i},n_{i}\mathrm{mod}N_{0}}$, but when computing $\eta(\bar{a},\bar{n})$ or $\gamma(\bar{a},\bar{n})$ it is essential that $n_{i}\in\mathbb{Z}$.
@@
The result about explicit generators of ¯$\bar{\mathcal{L}}^{[v_{0}]}$ can be stated as follows.
Let ¯$\bar{\mathcal{F}}^{0}_{\gamma,-N}$ be the image of $\mathcal{F}^{0}_{\gamma,-N}$ in ¯$\bar{\mathcal{L}}_{W(k)}$.
Let $\iota=p^{m}(qp^{M-1}\alpha-(q-1)\beta)\in\mathfrak{A}_{1}^{+}(p)$ be the standard presentation from Section 2.2. Then $|\iota|\leqslant p^{M-1}(p-1)b^{*}$, $p^{m}\alpha$ and $p^{m}\beta/r^{*}$ are integers not both divisible by $p$. We have also that $p^{m}(q-1)\beta/b^{*}=p^{m+M-1-A}$ where $10\leqslant A\leqslant M-$ 1, compare to Prop. 2.4b). So, by setting $\gamma(\iota):=p^{A}\alpha$ and $M(\iota):=m-A$ we obtain the following presentation of $\iota$:
@@
The condition $\iota\leqslant p^{M-1}(p-1)b^{*}$ holds automatically when we work with elements of $\bar{\mathcal{N}}^{{\dagger}}[0]$. Indeed, if $t^{-\iota}[D_{\bar{a}\bar{n}}]\in\bar{\mathcal{N}}^{{\dagger}}[0]$ then $-\iota\geqslant-qp^{M-1}|\bar{a}|$. So, if $\iota>p^{M-1}(p-1)b^{*}$ then $|\bar{a}|>(p-1)b^{*}/q$. By condition C3 from Section 2.1 it is equivalent to $|\bar{a}|\geqslant(p-1)v_{0}=v_{0}U(p-1.0)$ and then $[D_{\bar{a}\bar{n}}]\in\bar{\mathcal{L}}(p)_{W(k)}=0$.
For any $\iota\in\mathfrak{A}^{+}_{1}(p)$, fix a choice of $m_{\iota}\geqslant r(\iota)$, where $r(\iota)$ is the maximal integer such that $\iota p^{r(\iota)}\leqslant p^{M-1}(p-1)b^{*}$.
@@
¯$\bar{\mathcal{L}}^{[v_{0}]}$ is the minimal ideal in $\bar{\mathcal{L}}$ such that for all $\iota\in\mathfrak{A}^{+}_{1}(p)$ given with above notation (4.1), ¯$\bar{\mathcal{F}}^{0}_{\gamma(\iota),-(M(\iota)+m_{\iota})}\in\bar{\mathcal{L}}^{[v_{0}]}_{W(k)}$.
@@
All ¯$\bar{\mathcal{F}}_{\gamma,-N}$ appear as linear combinations of the Lie monomials $[D_{\bar{a}\bar{n}}]$ such that $\gamma=\gamma(\bar{a},\bar{n})=a_{1}p^{n_{1}}+\ldots+a_{s}p^{n_{s}}$ with $a_{i}\in\mathbb{Z}^{0}(p)$. There is a lot of monomials containing the generators $D_{0n}$ which do not contribute essentially to the condition $\gamma=\gamma(\bar{a},\bar{n})$. It would be natural to switch from the Lie algebra $\mathcal{L}$ to its quotient $\mathcal{L}^{o}=\mathcal{L}/(D_{0})$ which is related just to the generators $\{D_{an}\ |\ a\in\mathbb{Z}^{+}(p)\}$. Then we obtain the description of the corresponding ramification ideal $\mathcal{L}^{o(v_{0})}$ in terms of generators which all contribute (essentially) to the condition $\gamma=\gamma(\bar{a},\bar{n})$. As a matter of fact, this means that we fixed a lift $\hat{\sigma}$ of the Frobenius automorphism to $\mathcal{G}_{<p,M}$ and work with the Galois group of the totally ramified field extension $\mathcal{K}_{<p,M}^{\langle\hat{\sigma}\rangle}$. This field appears very naturally in the traditional local class field theory as well as in its nilpotent version [24].
@@ 4.4. A recurrent relation
Here $\bar{x}_{1}\in\bar{\mathcal{N}}^{{\dagger}}[0,1]$ and $V_{\iota}=\bar{\mathcal{V}}^{0}B^{{\dagger}}(D_{\iota 0}^{{\dagger}})$. We know that all $\sigma^{n}V_{\iota}$ generate the algebra ¯$\bar{\mathcal{L}}^{[v_{0}]}_{W(k)}$. Clearly, this property is not affected if we replace them modulo the ideal [¯$[\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}]_{W(k)}$.
Let $E_{0}=\exp(\sigma^{M-1}\bar{e}_{1}^{(q)})$ and $\exp(\Omega(\sigma^{M-1}\bar{e}_{1}^{(q)}))=E_{0}+\mathcal{E}_{0}$. In particular, $E_{0}=\sum_{a\in\mathbb{Z}^{0}(p)}\sigma^{M-1}\eta(\bar{a})t_{1}^{-qp^{M-1}|\bar{a}|}D_{\bar{a},M-1}$ and
@@
Under the agreement from Section 4.1 we have the following property:
@@
$\exp(l+l^{\prime})\cdot\exp(-l)-1=l^{\prime\prime}\in\bar{\mathcal{N}}^{{\dagger}}[0,1]\,,$
and $l^{\prime}\equiv l^{\prime\prime}\,\mathrm{mod}\,[l^{\prime},\bar{\mathcal{L}}]_{W(k)}$.
Proof.
This follows from a well-known formula for $\exp(-X)\exp(X+\delta)$ modulo the terms where $\delta$ appears at least twice, cf. e.g. [15], Theorem 4.22. ∎
Applying exponential to (4.2) and using $\sum_{\iota}t_{1}^{-\iota}V_{\iota}\in\bar{\mathcal{N}}^{{\dagger}}[0,1]$ we obtain
$E_{0}+\mathcal{E}_{0}\equiv(1+\sigma\bar{x}_{1})E_{0}(1-\bar{x}_{1})\left(1+\sum_{\iota}t_{1}^{-\iota}V_{\iota}\right)\,\mathrm{mod}\,[\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}]_{O^{\prime}_{M}(\mathcal{K}^{\prime})}\,.$
Then using that $\bar{x}_{1},\sigma\bar{x}_{1}\in\bar{\mathcal{N}}^{{\dagger}}[0,1]$ we rewrite this relation as follows:
Note also that by above Lemma 4.2 this relation is still in ¯$\bar{\mathcal{L}}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$.
@@
Let $\iota\in\mathfrak{A}^{0}_{1}$. Set (as earlier in the context of $\iota\in\mathfrak{A}_{1}^{+}(p)$)
—  $r(\iota):=\max\{r\ |\ \iota p^{r}\in\mathfrak{A}_{1}^{0}\}=\max\{r\ |\ |\iota p^{r}|\leqslant p^{M-1}(p-1)b^{*}\}$.
$m_{\iota}$ to be some fixed integer $\geqslant r(\iota)$.
@@
where the last sum is taken over $\iota\in\mathfrak{A}^{\mathrm{prim}}_{1}$, $r\geqslant 0$ and all ¯$\bar{\mathcal{F}}[\iota p^{r}]\in\bar{\mathcal{A}}_{W(k)}$.
Remark.
The elements ¯$\bar{\mathcal{F}}[\iota p^{r}]$ play a crucial role in the proof of Theorem 4.1. Namely, Proposition 4.3b) below establishes their relation with elements $V_{\iota}$ (which generate the ideal ¯$\bar{\mathcal{L}}^{[v_{0}]}$ ) and proposition 4.5 below expresses ¯$\bar{\mathcal{F}}[\iota p^{r}]$ in the form of Lie elements ¯$\bar{\mathcal{F}}^{0}_{\gamma(\iota),-M(\iota)+m_{\iota}}$. We are going to use now that ¯$\bar{\mathcal{F}}[\iota p^{r}]\in\bar{\mathcal{L}}_{W(k)}$ because the proof of Proposition 4.5 from Sections 4.6-4.7 does not rely on any constructions from this Section.
Let ¯$\bar{\mathcal{F}}[\iota p^{r}]_{0}:=\bar{\mathcal{F}}[\iota p^{r}]-\sigma\bar{\mathcal{F}}[\iota p^{r-1}]\in\bar{\mathcal{L}}_{W_{M}(k)}$. Then
@@
a) $\bar{x}_{1}\equiv-\sum_{\iota,r}\bar{\mathcal{F}}[\iota p^{r}]t_{1}^{-\iota p^{r}}\,\mathrm{mod}\,\bar{\mathcal{L}}^{[v_{0}]}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}+\bar{\mathcal{N}}^{{\dagger}}(p-1)$, where the sum is taken over all $\iota\in\mathfrak{A}^{\mathrm{prim}}_{1}$ and $0\leqslant r\leqslant m_{\iota}$;
b) if $\iota\in\mathfrak{A}_{1}^{+}(p)$, then $V_{\iota}\equiv\sigma^{-m_{\iota}}\bar{\mathcal{F}}[\iota p^{m_{\iota}}]\,\mathrm{mod}\,\left[\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}\right]_{W(k)}$.
@@
Apply induction on $1\leqslant w<p$ to prove the proposition modulo commutators $C_{w}(\bar{\mathcal{L}})_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$ for all $1\leqslant w\leqslant p$. Clearly, a) holds with $w=1$.
@@
(4.4) $\sum_{\mathrm{ch}\iota=1}t_{1}^{-\iota}V_{\iota}+\sigma{\bar{x}}_{1}^{1}-\bar{x}_{1}^{1}\equiv E_{0}^{-1}\mathcal{E}_{0}-\left(E_{0}^{-1}\sigma\bar{x}_{1}^{1}E_{0}-\sigma\bar{x}^{1}_{1}\right)\equiv$
$\sum_{r\leqslant m_{\iota}+1,\iota}\bar{\mathcal{F}}[\iota p^{r}]_{0}t_{1}^{-\iota p^{r}}\,\mathrm{mod}\,\left([\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}]+C_{w+1}(\bar{\mathcal{L}})\right)_{O^{\prime}_{M}(\mathcal{K}^{\prime})}+\bar{\mathcal{N}}^{{\dagger}}(p-1)\,.$
Since the LHS belongs to $\bar{\mathcal{N}}^{{\dagger}}[0,1]$ the above congruence implies that for $r(i)<r\leqslant m_{\iota}+1$, ¯$\bar{\mathcal{F}}[\iota p^{r}]_{0}\in\left([\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}]+C_{w+1}(\bar{\mathcal{L}})\right)_{W(k)}$. (Indeed, if $t_{1}^{-\iota p^{r}}[D_{\bar{a}\bar{n}}]\in\bar{\mathcal{N}}[0,1]$ then $qp^{M-1}|\bar{a}|\geqslant\iota p^{r}$ implies $|\bar{a}|>(p-1)v_{0}$ and $[D_{\bar{a}\bar{n}}]\in\bar{\mathcal{L}}(p)_{W(k)}=0$.)
@@
If $\iota<0$ then the expression for the appropriate part of $\bar{x}_{1}$ will be provided modulo ([¯$\left([\bar{\mathcal{L}}^{[v_{0}]},\bar{\mathcal{L}}]+C_{w+1}(\bar{\mathcal{L}})\right)_{O^{\prime}_{M}(\mathcal{K}^{\prime})}+\bar{\mathcal{N}}^{{\dagger}}(p-1)$ with
$\mathcal{S}\left(\sum_{r}\bar{\mathcal{F}}[\iota p^{r}]_{0}t_{1}^{-\iota p^{r}}\right)=-\sum_{j\geqslant 0}\sigma^{j}\left(\sum_{\iota,r}\bar{\mathcal{F}}[\iota p^{r}]_{0}t_{1}^{-\iota p^{r}}\right)=-\sum_{\iota,r\leqslant m_{\iota}}\bar{\mathcal{F}}[\iota p^{r}]t^{-\iota p^{r}}\,.$
Note that the elements from $\bar{\mathcal{N}}^{{\dagger}}[0](p-1)\subset\bar{\mathcal{L}}_{\mathrm{m}^{\prime}}$ will not contribute to $\sum_{\iota\in\mathfrak{A}_{1}^{+}(p)}t^{-\iota}V_{\iota}$. This proves part a) in the case $\iota<0$.
@@
$\mathcal{R}\left(\sum_{0\leqslant r\leqslant m_{\iota}}\bar{\mathcal{F}}[\iota p^{r}]_{0}t_{1}^{-\iota p^{r}}\right)\equiv\sum_{0\leqslant r\leqslant m_{\iota}}\sigma^{-r}\bar{\mathcal{F}}[\iota p^{r}]_{0}t_{1}^{-\iota}\equiv-\sigma^{-m_{\iota}}\bar{\mathcal{F}}[\iota p^{m_{\iota}}]t_{1}^{-\iota}\,.$
This proves part b) modulo $C_{w+1}(\bar{\mathcal{L}})_{W(k)}$.
@@
$\mathcal{S}\left(\sum_{0\leqslant r\leqslant m_{\iota}}\bar{\mathcal{F}}[\iota p^{r}]_{0}t_{1}^{-\iota p^{r}}\right)=\sum_{0\leqslant r\leqslant m_{\iota}}\sum_{0\leqslant r_{1}<r}\sigma^{r_{1}}\left(t_{1}^{-\iota}\sigma^{-r}\bar{\mathcal{F}}[\iota p^{r}]_{0}\right)$
$\equiv-\sum_{0\leqslant r\leqslant m_{\iota}}\sum_{r\leqslant r_{1}\leqslant m_{\iota}}\sigma^{r_{1}}\left(t_{1}^{-\iota}\sigma^{-r}\bar{\mathcal{F}}[\iota p^{r}]_{0}\right)$
$\equiv-\sum_{0\leqslant r_{1}\leqslant m_{\iota}}t_{1}^{-\iota p^{r_{1}}}\sigma^{r_{1}}\left(\sum_{r_{1}\leqslant r\leqslant m_{\iota}}\sigma^{-r}\bar{\mathcal{F}}[\iota p^{r}]_{0}\right)\equiv$
$\equiv-\sum_{0\leqslant r_{1}\leqslant m_{\iota}}t_{1}^{-\iota p^{r_{1}}}(\bar{\mathcal{F}}[\iota p^{r_{1}}]_{0}+\sigma^{-1}\bar{\mathcal{F}}[\iota p^{r_{1}+1}]_{0}+\ldots+\sigma^{r_{1}-m_{\iota}}\bar{\mathcal{F}}[\iota p^{m_{\iota}}]_{0})\equiv$
$-\sum_{0\leqslant r\leqslant m_{\iota}}\bar{\mathcal{F}}[\iota p^{r}]t_{1}^{-\iota p^{r}}$
@@
$\sum_{0\leqslant r_{1},r\leqslant m_{\iota}}\sigma^{r_{1}}\left(t_{1}^{-\iota}\sigma^{-r}\bar{\mathcal{F}}[\iota p^{r}]_{0}\right)=\sum_{r_{1}}t_{1}^{-\iota p^{r_{1}}}\sigma^{r_{1}}\left(\sum_{0\leqslant r\leqslant m_{\iota}}\sigma^{-r}\bar{\mathcal{F}}[\iota p^{r}]_{0}\right)\in\bar{\mathcal{L}}^{[v_{0}]}_{O^{\prime}_{M}(\mathcal{K}^{\prime})}$